### v6

@@ 2.5 Tightly focused light in highly anisotropic materials
In the extreme case of strong linear interaction between highly anisotropic materials and tightly focused light fields — an area at the forefront of both industry and academia in laser processing, aberration correction, and inverse focal engineering — our LFCO model provides a unified solution, as shown in Fig.˜8. Here, we show the underlying mechanism of Raman spike[161] (Fig.˜8b) in CR, the reverse engineering technique for the targeted vector complex field at the focal plane free of Zernike polynomial (Fig.˜8c1), the “Cherenkov cone” outside the inner beam (Fig.˜8c2) marking the computational boundary, the R-L and o-e decomposition analysis for high-N.A. pump (Fig.˜8c4), and double conical refraction and the optical field self-twisting effects it induces (Fig.˜8d1).
@@
Using the crytal-version vector angular spectrum method(ASM), i.e. this LFCO model, without any optimization, we explored various tightly focused field distributions while operating at the edge of the sampling theorem, with each subfigure (except Fig.˜8c2) in Fig.˜8 nearing the computational limits just before aliasing errors occur. Initially, we believed that the origin of the outer cone in Fig.˜8c2 was physical because: (1) it began forming even before propagation, and (2) it appeared at the center of the image rather than being reflected from the edges. However, we later discovered that it was still caused by aliasing errors originating from circular convolution. This conclusion was drawn because, in order to shift the focal plane of Fig.˜8c2 inside the LN materiral, we set the incident field distribution at the material’s entrance to a LP Gaussian beam with a N.A. close to 1 (see the top row of Fig.˜8c3), propagating backward by mm. However, when was set to zero, the outer cone disappeared. In contrast, Fig.˜8c4 uses the transfer function of the objective lens instead of backward propagation to shift the focal plane below the upper surface of LN, completely eliminating the aliasing error.
We find that by setting negative, thereby transforming the refractive index surface from an ellipsoid to a dielectric-type two-sheet hyperboloid[114, 118, 119, 113, 122], the two optical axes can be brought closer together, as shown in Fig.˜8d2. For KTP crystal at 1064 nm, the angle between the two optical axes decreases from 34.6 in the ellipsoidal case to 7.7 in the hyperboloidal case, allowing us to compute the full evolution inside and outside the crystal in a non-paraxial manner over relatively longer propagation distances of the entire crystal-2f system. Optical field knotting in the 1st and 3rd column of Fig.˜8d1 are beyond our present mathematical understanding and techniques for explanation.
@@ 2.6 Superstructure of this LCO model: NCO QNCO
Both the NFs and the pumps involved in the interaction, are constrained by their own monochromatic passive LCO wave equation, necessitating independent diffraction as the crystal’s eigenmodes (Fig.˜5d,e,g,h).
The NCO parametric frequency conversion process must first satisfy energy conservation, followed by momentum conservation (often described as wave vector or phase matching), which further tests the precision of LCO eigenvalue calculations (Fig.˜5d,e). If the anisotropy of the second-order nonlinear coefficient tensor is additionally involved, the accurate computation of eigenvectors (Fig.˜5g,h) in the frame[277, 13, 3, 4] within LCO must also be ensured as a prerequisite.
@@
Having conducted large-scale numerical validation of known NCO phenomena in Section˜2.2, thereby indirectly substantiating the LCO model advanced in this paper, it is imperative to further establish the model’s forward-looking predictive capabilities, for the potentially vast range of yet-unknown NCO phenomena.
To this purpose, we propose simulations of spontaneous parametric down-conversion(SPDC) in 1D, 2D, and 3D NPCs in Fig.˜9 (and Fig. SB6), as a preliminary step toward developing a comprehensive theoretical and mathematical FO framework for quantum NCO(QNCO) nonlinear quantum optics(NQO) QO.
@@ 3 Discussion
Conversely, numerically stable models that allow for the treatment of thick non-Hermitian slabs fail entirely to resolve EPs. These models share the same structure of classical plane wave , whose two distinct forms based on real spherical/rectangular coordinates, with different independent variables , both satisfy the same wave/characteristic Eq.˜2. All models choosing as input variable typically requires numerical solutions for quartics, while the rest models depends on possess simple closed-form solutions to biquadratics. However, the former satisfies FT and boundary conditions, while the latter does not.
@@ 4.1 Boundary conditions for laboratory settings
Then the wave equation (see Supplementary Note 2), together with Eq.˜6, contains all the necessary information for solving the distribution of electromagnetic field inside and outside the slab made of typical optical materials, provided that a two-dimensional(2D) distribution of the vector pump (’s transverse components ) in front of the slab () is given, which means neither divergence equations[191, 299, 231, 240, 91, 90, 300, 301] nor boundary conditions[191, 299] are even required.
@@
The trade-off is that, for each incident field, either two transmission fields passing through the anti-reflective(AR) coating/nanostructure or two reflected fields bounced back[304] by the high-reflective(HR) coating can be calculated relatively accurately, while finer effects such as the photon spin-Hall effect[306, 307] and the additional lateral shift[243] cannot be revealed by this model.
@@ 4.2 Eigensystem corrections for phase continuity
where both the input real ray direction for uniform eigenvalue and the output non-uniform eigenvalue are ultimately functions of , and mutually coupled, as illustrated in Fig.˜10a2.
One can unravel this coupling by first discarding the imaginary part of the complex non-uniform wave vector in Eq.˜8b, retaining only the real part ’s unit vector from Eq.˜9a as the ray direction in space. The input variable is then inserted into Eq.˜9b to form a transcendental equation which asymptotically converges to the ground true eigenvalue of the booker quartic on condition that
where material absorption is sufficiently low with typically when a transmission spectrum is still present, as exemplified by Brenier’s laser crystal[139, 81] in Fig.˜7b,e.
Berry & Dennis’s input variable can be obtained from Fig.˜10a2 as an intermediate data. Inserting it into Fig.˜10b yields Berry & Dennis’s eigensystem , where the eigenvector also converges to the ground truth under the condition of Eq.˜10 (see Supplementary Note 2.2), after undergoing the operation shown in Fig.˜10c.
@@ 4.3 Overall flowchart of this LFCO model
Tables˜1 and 2 briefly presents the categorical meanings of the graphical elements (nodes and arrows) as ‘CLASSes’ in Figs.˜6, 10 and 11. The physical/mathematical significance of the specific instantiated ‘OBJECTs’ within each ‘CLASS’ can be found in Supplementary Notes 2-5.
Starting from the atomic inputs, we built the comprehensive workflow diagram in Fig.˜11 from the bottom up, referencing three submodules from Fig.˜10a,c. All elements in the flowcharts (Figs.˜6, 10 and 11) of this article are accompanied by step-by-step derivations and detailed explanations provided in the Supplementary Notes.
Supplementary information
Acknowledgements
@@
Declarations
Funding This work was supported by the Natural Science Foundation of Jiangsu Province (BK20240005) and the Fundamental Research Funds for the Central Universities (021314380268).
Competing interests The authors declare no competing interests.
Data availability All data generated by this trilogy, along with the parameter file for reconstructing any (sub)figure, can be requested from the first author.
Code availability The trilogy, including scalar/vector L/NFCO, as an integrated entity, will be fully open-sourced upon the completion of the trilogy.
Author contribution C.X. conceived, led, and built the entire trilogy starting from empty-nothing (down to atom, up to soul) independently, with his unique long-chain/tree/web (non)linear logic, interspersed with leap-like flashes of divine apocalypto, right before the AI era and his own liberation from VPN restrictions. Y.Z. warmly approved, consistently supported, and responsibly supervised the whole trilogy.

### v7

@@ 2.5 Tightly focused light in highly anisotropic materials
In the extreme case of strong linear interaction between highly anisotropic materials and tightly focused light fields — an area at the forefront of both industry and academia in laser processing, aberration correction, and inverse focal engineering — our LFCO model provides a unified solution, as shown in Fig. 8. Here, we show the underlying mechanism of Raman spike[161] (Fig. 8b) in CR, the reverse engineering technique for the targeted vector complex field at the focal plane free of Zernike polynomial (Fig. 8c1), the “Cherenkov cone” outside the inner beam (Fig. 8c2) marking the computational boundary, the R-L and o-e decomposition analysis for high-N.A. pump (Fig. 8c4), and double conical refraction and the optical field self-twisting effects it induces (Fig. 8d1).
@@
Using the crytal-version vector angular spectrum method(ASM), i.e. this LFCO model, without any optimization, we explored various tightly focused field distributions while operating at the edge of the sampling theorem, with each subfigure (except Fig. 8c2) in Fig. 8 nearing the computational limits just before aliasing errors occur. Initially, we believed that the origin of the outer cone in Fig. 8c2 was physical because: (1) it began forming even before propagation, and (2) it appeared at the center of the image rather than being reflected from the edges. However, we later discovered that it was still caused by aliasing errors originating from circular convolution. This conclusion was drawn because, in order to shift the focal plane of Fig. 8c2 inside the LN materiral, we set the incident field distribution at the material’s entrance to a LP Gaussian beam with a N.A. close to 1 (see the top row of Fig. 8c3), propagating backward by mm. However, when was set to zero, the outer cone disappeared. In contrast, Fig. 8c4 uses the transfer function of the objective lens instead of backward propagation to shift the focal plane below the upper surface of LN, completely eliminating the aliasing error.
We find that by setting negative, thereby transforming the refractive index surface from an ellipsoid to a dielectric-type two-sheet hyperboloid[114, 118, 119, 113, 122], the two optical axes can be brought closer together, as shown in Fig. 8d2. For KTP crystal at 1064 nm, the angle between the two optical axes decreases from 34.6 in the ellipsoidal case to 7.7 in the hyperboloidal case, allowing us to compute the full evolution inside and outside the crystal in a non-paraxial manner over relatively longer propagation distances of the entire crystal-2f system. Optical field knotting in the 1st and 3rd column of Fig. 8d1 are beyond our present mathematical understanding and techniques for explanation.
@@ 2.6 Superstructure of this LCO model: NCO QNCO
Both the NFs and the pumps involved in the interaction, are constrained by their own monochromatic passive LCO wave equation, necessitating independent diffraction as the crystal’s eigenmodes (Fig. 5d,e,g,h).
The NCO parametric frequency conversion process must first satisfy energy conservation, followed by momentum conservation (often described as wave vector or phase matching), which further tests the precision of LCO eigenvalue calculations (Fig. 5d,e). If the anisotropy of the second-order nonlinear coefficient tensor is additionally involved, the accurate computation of eigenvectors (Fig. 5g,h) in the frame[277, 13, 3, 4] within LCO must also be ensured as a prerequisite.
@@
Having conducted large-scale numerical validation of known NCO phenomena in Section 2.2, thereby indirectly substantiating the LCO model advanced in this paper, it is imperative to further establish the model’s forward-looking predictive capabilities, for the potentially vast range of yet-unknown NCO phenomena.
To this purpose, we propose simulations of spontaneous parametric down-conversion(SPDC) in 1D, 2D, and 3D NPCs in Fig. 9 (and Fig. SB6), as a preliminary step toward developing a comprehensive theoretical and mathematical FO framework for quantum NCO(QNCO) nonlinear quantum optics(NQO) QO.
@@ 3 Discussion
Conversely, numerically stable models that allow for the treatment of thick non-Hermitian slabs fail entirely to resolve EPs. These models share the same structure of classical plane wave , whose two distinct forms based on real spherical/rectangular coordinates, with different independent variables , both satisfy the same wave/characteristic Eq. 2. All models choosing as input variable typically requires numerical solutions for quartics, while the rest models depends on possess simple closed-form solutions to biquadratics. However, the former satisfies FT and boundary conditions, while the latter does not.
@@ 4.1 Boundary conditions for laboratory settings
Then the wave equation (see Supplementary Note 2), together with Eq. 6, contains all the necessary information for solving the distribution of electromagnetic field inside and outside the slab made of typical optical materials, provided that a two-dimensional(2D) distribution of the vector pump (’s transverse components ) in front of the slab () is given, which means neither divergence equations[191, 292, 231, 240, 91, 90, 293, 294, 74, 75, 76, 77, 78, 295] nor boundary conditions[191, 292] are even required.
@@
The trade-off is that, for each incident field, either two transmission fields passing through the anti-reflective(AR) coating/nanostructure or two reflected fields bounced back[298, 299] by the high-reflective(HR) coating can be calculated relatively accurately, while finer effects such as the photon spin-Hall effect[300, 301] and the additional lateral shift[243] cannot be revealed by this model.
@@ 4.2 Eigensystem corrections for phase continuity
where both the input real ray direction for uniform eigenvalue and the output non-uniform eigenvalue are ultimately functions of , and mutually coupled, as illustrated in Fig. 10a2.
One can unravel this coupling by first discarding the imaginary part of the complex non-uniform wave vector in Eq. 8b, retaining only the real part ’s unit vector from Eq. 9a as the ray direction in space. The input variable is then inserted into Eq. 9b to form a transcendental equation which asymptotically converges to the ground true eigenvalue of the booker quartic on condition that
where material absorption is sufficiently low with typically when a transmission spectrum is still present, as exemplified by Brenier’s laser crystal[139, 81] in Fig. 7b,e.
Berry & Dennis’s input variable can be obtained from Fig. 10a2 as an intermediate data. Inserting it into Fig. 10b yields Berry & Dennis’s eigensystem , where the eigenvector also converges to the ground truth under the condition of Eq. 10 (see Supplementary Note 2.2), after undergoing the operation shown in Fig. 10c.
@@ 4.3 Overall flowchart of this LFCO model
Starting from the atomic inputs, we built the comprehensive workflow diagram in Fig. 11 from the bottom up, referencing three submodules from Fig. 10a,c. All elements in the flowcharts (Figs. 6, 10 and 11) of this article are accompanied by step-by-step derivations and detailed explanations provided in the Supplementary Notes.
Tables 1 and 2 briefly presents the categorical meanings of the graphical elements (nodes and arrows) as ‘CLASSes’ in Figs. 6, 10 and 11. The physical/mathematical significance of the specific instantiated ‘OBJECTs’ within each ‘CLASS’ can be found in Supplementary Notes 2-5.
@@
This work was supported by the Natural Science Foundation of Jiangsu Province (BK20240005) and the Fundamental Research Funds for the Central Universities (021314380268).
C.X. conceived, led, and built the entire trilogy starting from empty-nothing (down to atom, up to soul) independently, with his unique long-chain/tree/web (non)linear logic, interspersed with leap-like flashes of divine apocalypto, right before the AI era and his own liberation from VPN restrictions. Y.Z. warmly approved, consistently supported, and responsibly supervised the whole trilogy.
All data generated by this trilogy, along with the parameter file for reconstructing any (sub)figure, can be requested from the first author.